OpenStudy (anonymous):

One mole of N2 gas at STP will occupy what volume? 44.8 L 22.4 L 44.2 L 88.4 L none of the above How in the world do I figure this out?

OpenStudy (anonymous):

textbook

OpenStudy (ash2326):

one mole of any gas will occupy 22.4 liters

OpenStudy (ash2326):

at STP

OpenStudy (anonymous):

If I could find it in my textbook I would have. OS is my last resort. I don't expect people to do my homework @panlac01

OpenStudy (anonymous):

I am still looking for a textbook for quick and dirty tips on chemistry to upload on my website. I think not until this fall when I get an access to mastering courses

OpenStudy (anonymous):

lol

OpenStudy (ash2326):

@rebeccaskell94 Please check this, you'll understand how it's 22.4 liters http://hyperphysics.phy-astr.gsu.edu/hbase/kinetic/idegas.html

OpenStudy (anonymous):

whoa o.o

OpenStudy (anonymous):

ay yai yai

OpenStudy (anonymous):

So If it is one molecule at N2 do I just multiply but 2 or?

OpenStudy (anonymous):

I don't think that is right. 28 is not an option

OpenStudy (anonymous):

@karatechopper are you good at this?

OpenStudy (karatechopper):

Beccs i am only starting bio :) far from chem

OpenStudy (anonymous):

darn lol anyone else you know that is good?

OpenStudy (hba):

easy one and u dont know how to do it

OpenStudy (anonymous):

I need help not patronization. Thanks

OpenStudy (hba):

ok ill help you out

OpenStudy (hba):

do u know the mole formula

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I haven't been studying chemistry for long

OpenStudy (hba):

Moles = Volume( in dm3) / 24

OpenStudy (hba):

no but damn it it is an ideal case

OpenStudy (hba):

"1 mole of an ideal gas at STP has volume = 22.4 L"

OpenStudy (anonymous):

oh

OpenStudy (hba):

^^ thats ur answer

OpenStudy (hba):

u wanna know how u get it

OpenStudy (hba):

PV = nRT;at STP:T = 273K , P = 1 atm V/n = RT/P = (0.08206 Latm/moleK) x (273K) / (1 atm) = 22.4 L/mole

OpenStudy (hba):

thats it i guess

OpenStudy (anonymous):

so...much. @ . @